OpenStudy (anonymous):

Directions: Simplify each expression by distributing and combining like termsi need hep with this 10-3(4z+1)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Ok... distributing means multiplying the "thing" in front of a parenthesis times the things inside the parenthesis So, for example, 2(x + 4).... distribute the 2.... multiply 2 times "x" PLUS 2 times 4 equals 2x + 2(4) = 2x + 8 In your problem, you have: 10-3(4z+1) so distribute the 3 across the parenthesis... 10 - 3(4z) - 3(1) The next step is to combine like terms... meaning collect the number terms together... all the "z" terms are already together, you just have to multiply 3(4z) out to get __z

OpenStudy (anonymous):

can u show me plz

OpenStudy (anonymous):

I did show you... the previous comment shows you exactly the first step which is distributing...

OpenStudy (anonymous):

how do u get the answers

OpenStudy (anonymous):

you take the starting expression, multiply it out like I just showed you... then there is one more step. You have to work with me here a little...

OpenStudy (anonymous):

ok what do i need to do i don't get it

OpenStudy (anonymous):

????

OpenStudy (anonymous):

I explained how to distribute... this is a pretty important concept. Can you please read through it again and try to see if you can understand it? I explained it in as simple terms as I could think of... But hey, one last try... "distribute" just means multiply... if you and your friend both have some money... the same amount, just to make things easy, and the magic wizard comes along and casts a spell that triples everything, then it looks like this... Starting: (your$ + your friend's $) After wizard: 3 x (your$ + yourfriend's$) = 3 x your$ + 3 x your friend's $ That is distributing... take the number in front of the parenthesis and multiply it across the 2 numbers in the parenthesis.... a(x + y) = ax + ay I don't know how else to explain it... I can't just give you an answer... you have to try to understand how to distribute, or you need to talk to your teacher about it.